OpenStudy (aaronq):

So i'm having some trouble, i'm supposed to figure out the temperature at which a protein denatures: \(protein_{native} \rightleftharpoons protein_{denatured}\) i have: \(\Delta G^o=-RTln(K_{eq*folding}\) \(\-321.2 J/mol=-(8.314 J/mol)(T)ln(K_{eq*folding}\) i know the protein will denature when \(K_{eq*folding\)<1 but i can't set it to 1, because ln(1)=0, if use ln(0.99) i get like 3800 K, which is nonsense. I tried taking the limit as \(K_{eq*folding\rightarrow 1-}\) but i get infinity. What do?

OpenStudy (abb0t):

Hmm... is it when K\(_{eq}\) < 1 or <0?

OpenStudy (aaronq):

\(K_{eq}<1\), because it's the reverse process

OpenStudy (abb0t):

I feel like I should know this, my last class senior year was a protein biochemistry lab. If i remember: \(\sf \color{red}{ [D_{eq}] + [N_{eq}]=1}\)

OpenStudy (aaronq):

hmm, that makes sense. but i don't know how i can use that

OpenStudy (aaronq):

i'm gonna graph it on excel and see what happens

OpenStudy (abb0t):

Yeah, show me the graph.

OpenStudy (aaronq):

okay i'll post it, give me 2 mins

OpenStudy (abb0t):

You can mayb see the point of fractional denaturation.

OpenStudy (abb0t):

\(\Delta\)G\(_o\) = -RT + ln\(\sf \color{purple}{\frac[D_{eq}}{1-[D]}\)

OpenStudy (abb0t):

ugh. stupid LaTex, Im trying to type it out beautifully. Lol

OpenStudy (aaronq):

haha i know it screws up on me too sometimes. So i just put all the constants together, and typed into wolfram because it's easier. http://www.wolframalpha.com/input/?i=T%3D38.63%2Fln%28K%29

OpenStudy (aaronq):

theres an asymptote at K=1

OpenStudy (abb0t):

Did you just use temp to denature?

OpenStudy (aaronq):

we'll i'm trying to find the temp at which it denatures. The initially gave me some data, and i found the thermodynamic parameters H, S, G for the folding process. And now, it's asking me to find the temp at which the protein denatures.

OpenStudy (aaronq):

so i figured that implies that Keq<1

OpenStudy (abb0t):

I would use \(\Delta\)G = \(\Delta\)H-T\(\Delta\)S in that case.

OpenStudy (aaronq):

so should use that and set \(\Delta G\)>0, for the folding reaction. Sorry, i'm tired lol

OpenStudy (abb0t):

It's ok. I should be studying right now for my microbiology exam tmrw morning but I'm on here \(\sf \color{pink}{:P}\) cuz now its bothering me that I don't rmr :/

OpenStudy (abb0t):

And yeah. Are you familiar with the van't hoff equation? I think this is similar.

OpenStudy (aaronq):

yeah, i was thnking about that, but, again, I'm not sure what i would set the ratio of the equilibrium constants to. Dude, i appreciate your help, but you should go study if you have an exam! lol i'll definitely msg you the answer (i'll ask my prof tmrw) if you're interested.

OpenStudy (aaronq):

haha i know! it gets pretty addictive. Nah man, you're a smart mofo, sometimes getting the answer is just not possible because there's not enough info n $h!t. But yeah, i'll let you know, peace !

OpenStudy (aaronq):

So, to find the denaturation temp, i had to use \(\Delta G=\Delta H-T\Delta S\) and at equilibrium \(\Delta G=0 \) and thats it lol

OpenStudy (aaronq):

@abb0t

OpenStudy (abb0t):

OMG, I was right. HAHA

OpenStudy (abb0t):

I think you could of used the slope also, which is something like \(\frac{\Delta S}{\Delta H}\) i think?